OpenStudy (anonymous):

Solve the given optimization problem by using substitution. Find the minimum value of f(x, y, z) = 2x^(2) + 2x + y^(2) − y + z^(2) − z − 5 subject to z = 2y. I actually tried to do this one and couldn't figure it out

OpenStudy (whpalmer4):

Hmm, don't you substitute \(2y\) for z throughout \(f(x,y,z)\) giving you\[f(x,y,z) =2x^2+2x+y^2-y+(2y)^2-(2y)-5 = 2x^2+2x+5y^2-3y-5\] Then take the partial derivatives of that with respect to \(x\) and \(y\) and set them equal to 0? Solve those two equations and back substitute to find \(z\). Isn't that how you work these?

OpenStudy (anonymous):

yeah but how do you find the minimum

OpenStudy (whpalmer4):

The point at which it is a minimum, or the value there?

OpenStudy (anonymous):

the value

OpenStudy (whpalmer4):

Just plug in the values. Do you have the values of \(x,y,z\) yet?

OpenStudy (anonymous):

i have x as -1/2 and y as 1/3

OpenStudy (anonymous):

y as 1/2

OpenStudy (whpalmer4):

x = -1/2, y = 3/10 are what I got. z = 2y = 6/10 = 3/5

OpenStudy (whpalmer4):

\[f(x,y,z) = 2x^2+2x+y^2-y+z^2-z-5\]Substitute \(z=2y\)\[f(x,y) = 2x^2+2x+y^2-y+(2y)^2-(2y)-5 \]\[f(x,y)= 2x^2+2x+5y^2-3y-5\] Agreed so far?

OpenStudy (anonymous):

yeah that makes sense so far

OpenStudy (whpalmer4):

\[\frac{\partial \left(f(x,y)\right)}{\partial x} = 4x+2\] \[\frac{\partial \left(f(x,y)\right)}{\partial y} = 10y-3\]

OpenStudy (whpalmer4):

Setting those equal to 0 and solving, we get \[4x+2=0\]\[4x=-2\]\[x=-\frac{1}{2}\] \[10y-3=0\]\[10y=3\]\[y=\frac{3}{10}\]\[z=2y\]\[z=2*\frac{3}{10} = \frac{3}{5}\] Agreed?

OpenStudy (anonymous):

okay yeah i got that, I'm still a little confused on how you got 5y^(2)

OpenStudy (whpalmer4):

Oh, remember \(z = 2y\) so \(z^2 = (2y)^2 = 4y^2\)

OpenStudy (anonymous):

i keep getting 3y^(2)

OpenStudy (whpalmer4):

and there's another \(y^2\) from the first part of the equation to add to that... I made the same mistake :-)

OpenStudy (anonymous):

wouldn't that make it 4y though? not 4y^(2)?

OpenStudy (whpalmer4):

if \(z = 2y\), \(z^2 = z*z = (2y)(2y) = 2*2*y*y = 4y^2\)

OpenStudy (anonymous):

ohhh gotcha haha my bad

OpenStudy (anonymous):

so then how do you find the minimum value?

OpenStudy (whpalmer4):

yeah, that's why I try to remember to put () around things when substituting...

OpenStudy (whpalmer4):

\[f(x,y,z) = 2x^2 + 2x+y^2-y+z^2-z-5\]Substitute in our values of \(x,y,z\) \[f(-\frac{1}{2},\frac{3}{10},\frac{6}{10}) = 2(-\frac{1}{2})^2 + 2(-\frac{1}{2}) + (\frac{3}{10})^2-(\frac{3}{10})+(\frac{6}{10})^2-(\frac{6}{10}) - 5\]\]\[=2(\frac{1}{4})-1+\frac{9}{100}-\frac{30}{100}+\frac{36}{100}-\frac{60}{100}-5\]\[=-6+\frac{1}{2}-\frac{45}{100} = -\frac{119}{20} = -5.95\]

OpenStudy (whpalmer4):

Here's a contour plot: You can see that \((-\frac{1}{2},\frac{3}{10})\) is right in the middle of the innermost contour line.

OpenStudy (anonymous):

oooh geez thats why i couldn't figure this one out but thank you so much!!! i really appreciate you showing the steps

OpenStudy (whpalmer4):

If you have the patience (or a programmable calculator), you can crank out some values nearby and verify that this is a minimum, not a maximum.

OpenStudy (whpalmer4):

Would you believe this is the first time I've solved one of these? :-)

OpenStudy (anonymous):

haha no way?!

OpenStudy (whpalmer4):

Here's a somewhat zoomed in contour plot, in color.

OpenStudy (anonymous):

well you did an awesome job :)

OpenStudy (anonymous):

thats the graph of it?

OpenStudy (anonymous):

and do you possibly have time to help me with one more? :) its a different type of problem i just need an example to see how to do it

OpenStudy (whpalmer4):

Gave me an excuse to experiment with Mathematica's contour plotting :-) Yes, lighter colors are larger values, darker colors are smaller. It's like a topographical map of the surface of the function — everything on one of those rings has the same value of \(f(x,y,z)\)

OpenStudy (whpalmer4):

Well, there's one way to find out! :-) Post the new question and tag me...

OpenStudy (anonymous):

how do i tag lol

OpenStudy (whpalmer4):

just put "@whpalmer4" in a post

OpenStudy (anonymous):

okay gotcha thank you so much! :)

OpenStudy (whpalmer4):

then I get a notification and a link I can click to go there without having to search for it.